If X and Y are independent Poisson variates with mean lambda 1 and lambda 2 respectively, find the probability that x+y=k

1
Expert's answer
2022-01-17T16:03:58-0500

Sums of independent Poisson random variables are Poisson random variables. Let "X" and "Y" be independent Poisson random variables with parameters "\u03bb_1" and "\u03bb_2," respectively. Define "\u03bb = \u03bb_1 + \u03bb_2" and "Z = X + Y." We claim that "Z" is a Poisson random variable with parameter "\u03bb."

Then


"P(X+Y=k)=P(Z=k)=\dfrac{e^{-\lambda}(\lambda)^k}{k!}"

"P(X+Y=k)=\dfrac{e^{-\lambda_1+\lambda_2}(\lambda_1+\lambda_2)^k}{k!}"